OpenStudy (anonymous):

@SolomonZelman I need to check my answer and I have two other parts that I need help with.

OpenStudy (solomonzelman):

Yes, how may I help you ?

OpenStudy (anonymous):

Iris has been studying an invasive population of snails. This particular snail has no local predators, so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences. 15 years when x is equal to the time unit. So 15 is equal to x. We don’t have P so let’s have 200 as P and let’s say that the rate doubles each year so we will have .5. P = amount you start with R = rate X = time f(x) = p(1 + r)^x f(15) = 200(1 + 0.5)^15 f(15) = 200(1.5)^15 f(15) = 200(437.893890380859) f(15) = 87578.7780761718 This will be the snail population in 15 years. The solution is irrational. A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population. g(x) = 200(1.03)^2x g(x) = 200(1.03^2)^x g(x) = 200(1.0609)^x g(x) = 212.18 This solution is irrational. Both solutions for the local and invasive are irrational.

OpenStudy (anonymous):

Are these correct?

OpenStudy (solomonzelman):

Lets start from here, P is the y-intercept R is the slope X (not the lower case x, but the one you wrote next to `P` and `R`) is the `f(x)` itself

OpenStudy (solomonzelman):

No

OpenStudy (solomonzelman):

X is the x.

OpenStudy (solomonzelman):

And r is the rate of an EXPONENTIAL FUNCTION.

OpenStudy (solomonzelman):

I know what you did incorrectly, I'll post the convertion

OpenStudy (anonymous):

Well it didnt give me P or R so I pluged in 200 for P and 0.5 for R and my tutor said that it was correct, I just wanted to check the second one.

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm g(x) = 200(1.03)^{2x} }\) \(\large\color{midnightblue}{ \rm g(x) = 200(1.03^2)^{x} }\) \(\large\color{midnightblue}{ \rm g(x) = 200(1.0609)^{x} }\) and where is 212.18 from ?

OpenStudy (anonymous):

200(1.0609) = 212.18 right?

OpenStudy (solomonzelman):

This is only when x=1

OpenStudy (anonymous):

so what is x?

OpenStudy (solomonzelman):

Assuming that each x unit is a year, this would be CORRECT, for ONE YEAR.

OpenStudy (anonymous):

Oh wait, it is 15 years for the first one so it has to be 15 years for this one as well

OpenStudy (solomonzelman):

It depends, on the question, \(\large\color{midnightblue}{ \rm g(x) = 200(1.0609)^{x} }\) (which is same as \(\large\color{midnightblue}{ \rm g(x) = 200(1.03)^{2x} }\) ) IS for how many years ?

OpenStudy (solomonzelman):

I would think that it is asking you for a 15-tear period.

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm g(x) = 200(1.0609)^{x} ~~⇒~~ g(x) = 200(1.0609)^{15}~~⇒~~? }\)

OpenStudy (solomonzelman):

you can find 15 by graphing x=15 and seeing where the curve and x=15 intersect.

OpenStudy (solomonzelman):

or just calculate it as I showed ... tell me if you need more help

OpenStudy (anonymous):

Iris wants to graph the invasive snail population to show the city council. Justify what the appropriate domain and range would be for the function f(x), what the y-intercept would be, and if the function is increasing or decreasing.

OpenStudy (anonymous):

heres the graph that I have @SolomonZelman

OpenStudy (solomonzelman):

how in the world do you get a line ?

OpenStudy (solomonzelman):

sorry about the world ;)

OpenStudy (anonymous):

i graphed the function and that is how it came up..

OpenStudy (anonymous):

Y = 200(1 + 0.5)^15

OpenStudy (anonymous):

It has a y-intercept of 87578.78 and a zero slope. The function is increasing because the snail population grows.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

I graphed the function in red x=15 in blue and the intersection point is labeled in black https://www.desmos.com/calculator/kztgaj4doi

OpenStudy (anonymous):

Thanks! :)

OpenStudy (solomonzelman):

Anytime !